OpenStudy (anonymous):

Find an equation of the tangent line to the curve at the given number. r=2+cos(theta) when theta=pi/4

OpenStudy (anonymous):

I have y=(2+cos(theta))(sin(theta)) x=(2+cos(theta))(cos(theta)) dy/dx=/\[((-\sin \theta)(\sin \theta)+(\cos \theta)(2+\cos \theta))\div((-\sin \theta)(\cos \theta)+(-\sin \theta)(2+\cos \theta))\] so at theta=pi/4 dy/dx=\[-\sqrt{2}/(\sqrt{2}+1)\]

OpenStudy (anonymous):

I set it up, y=mx+b \[\sqrt{2}+.5=((-\sqrt{2}/(\sqrt{2}+1))(\sqrt{2}+.5)+b\] just having issues finding b